# Accounting for multiplicity in machine learning benchmark performance

### v5

@@ 3.1 Notation
3.2 A coin-flip example
Multiplicity is, simply put, how the probability of an outcome changes when an experiment is performed multiple times.
Consider an experiment where a fair coin is flipped $n=20$ times and the number of successes is the number of times a classifier correctly predicts the outcome. Let the random variable $X$ denote the number of failures, $X\sim\mathcal{B}(n,1\text{-}\theta)$, where $\theta=0.5$. An unbiased estimator for $\theta$ is ${\hat{\theta}(X)=(n-X)/n}$. The corresponding estimate is referred to as accuracy. The probability of, say, $18$ correct predictions or more corresponds to $x=n-18$ failures or less, that is, the cumulative distribution $P(X\leq x|n,1\text{-}\theta)$. This corresponds to the cumulative distribution for the estimator $\hat{\theta}(X)=\frac{n\text{-}X}{n}$, and we have that $P(\hat{\theta}(X)\geq\frac{n\text{-}x}{n})=P(X\leq x)=0.00020$ for $\hat{\theta}\geq 0.90$. In other words, if $\theta=0.5$, there is a probability of $0.00020$ that the accuracy is $0.90$ or above in a series of $n=20$ trials. We can define the experiment described above as a Bernoulli trial, where success is defined as $X\leq x$, with corresponding probability $P(X\leq x)$, which we from now on write as $P_{x}$ for ease of notation.
If we have $m$ classifiers, we have a series of $m$ Bernoulli trials, and the number of successes follow the binomial distribution $\mathcal{B}(m,P_{x})$. Let $C_{x}$ be the random variable that denotes the number of successful experiments, that is, the number of classifiers that correctly predicts at least $n-x$ coin flips. Of special interest is the case where at least one experiment succeeds, $P(C_{x}>0)=1-P(C_{x}=0)$.
The true SOTA for coin-flip classifiers is $\theta_{SOTA}=0.5$, since it is impossible to train a classifier to predict coin flips. If $m=1,000$ classifiers independently give their predictions for $n=20$ flips, the probability of at least one classifier achieving an accuracy of at least $0.90$ is $0.181-P(C_{x}=0|m,P_{x})=$ 0.18. This means that using the top-ranked method as an estimate and ignoring multiplicity there is a substantial risk of grossly overestimating the SOTA.
The same mechanism of multiplicity applies when $m$ teams submit their classifier’s result on a test set, although in a more complex manner. We seek to describe the probability distribution for increasing levels of complexity, and by that offer a tool to avoid overestimation. We will in the following use the public competition as a general example, and refer to the teams’ classifiers simply as classifiers, for ease of notation.
3.3 The probability distribution of $\hat{\theta}^{\,\max}_{SOTA}(\mathbf{X})$
We define the SOTA probability of success, $\theta_{SOTA}$, as the maximum among the $m$ classifiers being evaluated, $\theta_{SOTA}=\max(\Theta)$, $\Theta=\{\theta_{1},\theta_{2},\ldots,\theta_{m}\}$. Oftentimes, we only have access to sample estimates, $\hat{\Theta}=\{\hat{\theta}_{1},\ldots,\hat{\theta}_{m}\}$. Let $\operatorname{\hat{\theta}^{\,max}_{SOTA}}(\mathbf{X})=\max(\{\hat{\theta}_{1}(X),\hat{\theta}_{2}(X),\ldots,\hat{\theta}_{m}(X)\})$ be the naive, sample maximum estimator for $\theta_{SOTA}$. In the coin flip classifier example above, we have that $\theta_{SOTA}=0.5$, but $P(\hat{\theta}^{\,\max}_{SOTA}(\mathbf{X})\geq 0.90)=0.18$, and it is clear that we risk an overoptimistic estimate.
Of interest is $P(\operatorname{\hat{\theta}^{\,max}_{SOTA}}(\mathbf{X})|m,n,\Theta)$, which we will denote by $P(\operatorname{\hat{\theta}^{\,max}_{SOTA}})$ for ease of notation. Similarly, we use $\operatorname{\hat{\theta}^{\,max}_{SOTA}}$ for $\operatorname{\hat{\theta}^{\,max}_{SOTA}}(\mathbf{X})$ when there is no risk of confusion. We here give $P(\operatorname{\hat{\theta}^{\,max}_{SOTA}})$ in the simple case of independent classifiers with identical probabilities of success: $\theta_{j}=\theta$ for all classifiers, $j=1,\ldots,m$. We discuss non-identical probabilities and dependency in Section 4.
Let $X$ be the random variable that denotes the number of failures (wrong predictions) for a given classifier, with $X\sim\mathcal{B}(n,1\text{-}\theta)$. The random sample, $\mathbf{X}=\{X_{1},X_{2},\ldots,X_{m}\}$ consists of the number of wrong predictions for each classifier $j=1,\ldots,m$ on a sample. Let $C_{x}$ be the random variable that denotes the number of classifiers with at most $x$ failures. We have that $C_{x}\sim\mathcal{B}(m,P_{x})$.
We are interested in the probability of at least one classifier having at most $x$ wrong predictions, because it corresponds to the probability function of the $\operatorname{\hat{\theta}^{\,max}_{SOTA}}$ estimator, $P(\operatorname{\hat{\theta}^{\,max}_{SOTA}}\geq\frac{n-x}{n})$. We can derive this from $P(C_{x}>0|m,n,\theta)$. This corresponds to the cdf for the random variable $Z$ denoting the most wrong predictions among all classifiers:
$F(z)$ $=P(C_{z}>0|m,n,\theta)=1-P(C_{z}=0|m,P_{z})$
$=1-(1-P_{z})^{m}.$ (1)
Let $f(z)=P(Z=z)$ denote the pmf of $Z$, the probability of at least one classifier having exactly $z$ wrong predictions, and no classifiers having fewer than $z$ wrong predictions. We then have
$f(z)$ $=F(z)-F(z-1)=1-(1-P_{z})^{m}-\{1-(1-P_{z-1})^{m}\}$
$=(1-P_{z-1})^{m}-(1-P_{z})^{m}=$
$=(1-\sum_{\ell=0}^{z-1}P(Z=\ell))^{m}-(1-\sum_{\ell=0}^{z-1}P(Z=\ell)-P(Z=z))^{m}$
$=-\sum_{k=1}^{m}(-1)^{k}\binom{m}{k}(1-\sum_{\ell=0}^{z-1}P(Z=\ell))^{m-k}P(Z=z)^{k}.$ (2)
The variable $Z$ corresponds to $\hat{\theta}_{SOTA}^{\,max}$ in the sense that the smallest number of failures among all classifiers corresponds to the sample maximum probability of success. Equations 3.3 and 3.3 can be used to analyse $\operatorname{\hat{\theta}^{\,max}_{SOTA}}$ for a given situation, as we do in the next section, where the impact that the three parameters have on the bias is also presented.
3.4 A simulated public competition example
The multiplicity mechanism that we described in the coin-flip example applies to public competitions where multiple teams submit the predictions made by their classifiers on a test set. The classifier with highest performance is declared winner, and is often referred to as SOTA later on. As exemplified in Section 3.2, multiplicity biases the naive sample maximum $\theta_{SOTA}$-estimator, and Section 3.3 gives us the tools to analyze its impact. In the following, we will use a significance level of $\alpha=0.05$ and two-sided CIs.
Consider a classification problem with a test set of size $n=3,000$, and a classifier with probability of success $\theta=0.90$. The pmf of a single $\hat{\theta}(X)$ is shown as the blue curve in Fig. 1, where the dotted line indicates $E(\hat{\theta})=\theta$. The 95% CI for $\hat{\theta}=0.90$ is $(0.8887,0.9105)$, and we denote it by $(\theta_{\alpha/2},\theta_{1\text{-}\alpha/2})$, displayed as a blue line on the horizontal axis.
If there are $m=1,000$ independent classifiers, each with probability of success $\theta=0.90$, the probability of at least one classifier achieving at least $\hat{\theta}=\theta_{1\text{-}\alpha/2}$ is close to 1. This comes as no surprise given the definition of a CI. It means that we can be pretty certain that the sample performance of the top ranked classifier is above the upper limit of the CI of the expected performance.
Consider a new classifier with probability of success $\theta_{m+1}=\theta_{1\text{-}\alpha/2}$. The corresponding pmf is displayed as the red curve in Fig. 1, with the dotted line indicating $E(\hat{\theta}_{m+1})=\theta_{1\text{-}\alpha/2}$. By common scientific standards, the expected performance is considered significantly better than the expected performance of any of the $1,000$ classifiers with $\theta=0.90$, and is of interest to the research community. However, the multiplicity effect can conceal this. In the competition context, this is not necessarily a problem, since each classifier has the same probability of achieving the highest sample performance given that they have the same $\theta$. In a scientific context, using $\operatorname{\hat{\theta}^{\,max}_{SOTA}}$ as an estimate for $\theta_{SOTA}$ without multiplicity corrections is problematic.
The upper limit of the $(1\text{-}\alpha)$ CI of $\operatorname{\hat{\theta}^{\,max}_{SOTA}}$ is $0.9213$, displayed on the horizontal axis in Fig. 1, denoted as $\theta^{m}_{1\text{-}\alpha/2}$. We have $E(\operatorname{\hat{\theta}^{\,max}_{SOTA}})=0.9173$, calculated from Eq. 3.3, also displayed on the horizontal axis in Fig. 1. The probability of the new classifier beating the upper limit of the sample maximum CI is very low: $P(\hat{\theta}_{m+1}\geq\theta^{m}_{1\text{-}\alpha/2})=0.02$, illustrated by the darker green area in Fig. 1. Similarly, $P(\hat{\theta}_{m+1}\geq E(\operatorname{\hat{\theta}^{\,max}_{SOTA}}))=0.10$ seen as the entire green area in Fig. 1. This means that by using the sample maximum as an estimator for SOTA, we can expect that 9 out of 10 classifiers that are significantly better will still be naively considered worse.
The pmfs of two single $\hat{\theta}(X)$ compared to $\operatorname{\hat{\theta}^{\,max}_{SOTA}}$ statistics
Figure 1: Against a mass of identical classifiers, a single significantly better classifier has little chance of beating the state of the art: The blue and the red curves show the pmfs of $\hat{\theta}(X)$ for classifiers with probability of success $\theta$ and $\theta_{1\text{-}\alpha/2}$, respectively. $\theta_{1\text{-}\alpha/2}$ is the upper limit of the CI of $\hat{\theta}(\mathbf{x})=\theta$. The CI is displayed as the blue horizontal line. $E(\operatorname{\hat{\theta}^{\,max}_{SOTA}})$ and $\theta^{m}_{1\text{-}\alpha/2}$ denote the expected value and the upper limit of the CI of the sample maximum estimator for $m=1,000$ classifiers. The shaded areas corresponds to $P(\hat{\theta}_{m+1}\geq\theta^{m}_{1\text{-}\alpha/2})$ and $P\big{(}\hat{\theta}_{m+1}\geq E(\operatorname{\hat{\theta}^{\,max}_{SOTA}})\big{)}$, and reflects the slim chances that the single better classifier has of beating the multiple worse classifiers.
For the entire distribution of $\operatorname{\hat{\theta}^{\,max}_{SOTA}}$, see Fig. 2.
Cdf and pmf for at least one classifier having at most $z$ failures, corresponding to the sample maximum estimator
(a) Cdf $F(z)$, see Eq. 3.3
(b) Pmf $f(z)$, see Eq. 3.3
Figure 2: The distribution and the probability function of the sample maximum estimator for $n=3,000$, $\theta=0.90$, and $m=1,000$. The horizontal axes denotes the number of failures, $z$, at the top and the corresponding sample maximum estimates, $\operatorname{\hat{\theta}^{\,max}_{SOTA}}$, at the bottom. Note that the horizontal axis is reversed compared to Fig. 1.
The distribution of $\operatorname{\hat{\theta}^{\,max}_{SOTA}}$ will vary with $m$, $n$ and $\theta$, and Table 1 shows the expectation and standard deviation for a small collection of examples. To complement the table, Fig. 3 a)-c) show the bias of $\operatorname{\hat{\theta}^{\,max}_{SOTA}}$ as a function of $m$, $n$ and $\theta$, respectively.
Expected values and standard deviations for $\operatorname{\hat{\theta}^{\,max}_{SOTA}}$
m n $\theta$ $E(\operatorname{\hat{\theta}^{\,max}_{SOTA}})$ $\sigma_{\operatorname{\hat{\theta}^{\,max}_{SOTA}}}$

### v6

@@ 5.2.1 Simulation of AUC
5.2.2 Simulation results for uncorrelated classifiers
To demonstrate the impact that multiplicity can have, we start with a simulated example similar to the one in Section 3.4. Consider a classification problem with $n=3,000$, $m=1,000$, where each classifier has an AUC of $a=0.90$, and the classifiers are independent.
Whereas the distribution of the sample maximum for probability of success is independent of class balance as long as the probability of success is independent of class, this is not true for the sample maximum of AUC, where the variance increases with class imbalance. We have chosen the same class balance as in the Melanoma challenge, $\pi=0.017$, as this reflects real world examples. For simplicity we choose $\mu_{-}=0$ and $\sigma_{-}=\sigma_{+}=1$. As we see from Eq. 10 the AUC is completely decided by the ratio between the difference in means and the pooled standard deviation. Changing the scale of one implies the same change of scale in the other. Perhaps the strongest assumption is that of equal variances, but we see unrealistically variable AUCs from letting the difference between the two be too large, so keeping them equal was the simplest choice. We then solve for $\mu_{+}$ numerically. Now we can generate scores from the resulting Normal distributions.
Figure 10(c) shows the simulated distribution of the sample maximum AUC over $10,000$ repetitions of a challenge where every classifier has AUC of $0.90$. The simulated expected value is $0.9562$, the standard deviation is $0.004459$, and the simulated 95% CI is $(0.9486,0.9662)$. Comparing this to the results in the upper row of Table 3 , we see that the bias is much larger for the sample maximum AUC. This is partly due to the large variance created by the class imbalance. Although the sample size is $n=3,000$, only $51$ of those belong to the positive class, creating large variance in the estimator.
5.2.3 Melanoma Classification
We will use the Melanoma Classification challenge from 2020 as an illustrative example of $AUC_{SOTA}$ estimation because it follows many of the common structures of public challenges.
The SIIM-ISIC Melanoma Classification (Zawacki et al., 2020) challenge consisted of more than $40,000$ images of skin lesions from various sites, split into training ($33,126$), validation ($3,295$) and test set ($7,687$). The data set is highly imbalanced, with less than $2\%$ melanomas, which reflects the reality in many cancer detection situations. A full description can be found in Rotemberg et al. (2021). The size of the test set is smaller than the recommendations of, for example, Roelofs et al. (2019), but reflects the reality of benchmark datasets and public challenges, as seen in, for example, the overview of Willemink et al. (2020). We want to emphasise that the shortcomings of this challenge are not unique, and that is precisely why it serves as a useful example.
The prize money was set to $10,000$ USD for the best-performing method, measured as AUC on the test set. Figure 11(a) shows the results. The team ‘All Data Are Ext’, consisting of three Kaggle Grandmasters, won the challenge, achieving an AUC of $0.9490$, and their method and strategy is explained in Ha et al. (2020). Although we show that the AUC is an optimistic estimate for the melanoma detection SOTA, we do not question their rank in the challenge. Quite the contrary; the winning team chose many of the verified strategies from Section 2. They used additional data from previous years’ challenges, and they also did data augmentation. They chose ensemble learning as a strategy to avoid overfitting, a strategy implicitly supported by Talagrand (1996). They also trained on multiple classes, a strategy that Feldman et al. (2019) showed will slow down the overfitting.
The observed maximum is an unlikely SOTA candidate, and to give a better estimate of the SOTA for the Melanoma challenge, we follow the same procedure as in Section 5.1: The challenge observations are shrunk in such a manner that simulating from the shrunk set, $AUC^{\prime}$, gives an expected sample maximum AUC of $0.9490$, the challenge sample maximum.
Use of shrinkage distribution to suggest a SOTA
(a) Shrinkage AUC distribution (black) based on the Melanoma challenge results (grey). The vertical red line shows the highest Melanoma challenge AUC of .949.949.949.949 and the horizontal red line shows a $95\%$ CI around this.
(b) A sample from the shrinkage distribution of (a). The vertical line is the expected sample maximum, engineered to coincide with the observed Melanoma challenge maximum of $0.949$. The horizontal line shows the $95\%$ CI of the sample maximum.
Figure 11: Histograms for AUC sample estimates shrunk toward $0.5$. In the shrinkage distribution the state of the art is an AUC of $0.8992$ which leads to the expected sample maximum AUC of 0.9490. AUCs below .6.6.6.6 not shown to make the more interesting upper region more readable.
Figure 11(a) shows the AUCs of the Melanoma challenge test sample shrunk toward $0.5$, with maximum $0.8992$. Figure 11(b) shows a corresponding sample, where the maximum is close to $0.9490$. There are 2013 teams with performance above the new SOTA candidate. The discrepancy between the sample maximum and the new SOTA candidate is big and has big implications. From an AUC close to $95\%$ to barely $90\%$ is a substantial drop in performance. Almost two thirds of the classifiers have performances above the new SOTA estimate.
As was shown in Section 4, illustrated in Figure 5(b), the bias is reduced for correlated classifiers, so the analysis assuming independence overestimates the bias. In this challenge, we can assume multiple sources of dependency: Access to the same training set, models pre-trained on the same data, similar models, etc. However, the wide CI for a single classifier tells the story of an estimator with large variance, and hence the bias of the sample maximum will nevertheless be substantial.
5.3 Future work
There are several interesting aspects that has not been addressed in this work, that the authors briefly touched upon while preparing the manuscript.
5.3.1 Correlation between classifiers in the AUC simulations
For the moment, we have not extended our AUC simulations to include correlations, but it should be possible to adopt a similar setup as in Sections 4.2 and 4.3: each classifier produces scores that correlate to those of a reference classifier. Let $S_{0^{+}}$ and $S_{0^{-}}$ be the scores of the reference classifier, for random variables from the positive and negative class, respectively.
and equivalently for the negative class. However, the joint distribution of $(S_{0},S_{j})$ is not uniquely defined by the correlation, and additional assumptions must therefore be made.
I
5.3.2 Approximate Bayesian Posterior for SOTA
In Section 5 we consider observed data, $\mathbf{x}$, as coming from a hierarchical model like that of Figure 1 and try to find a $P(\theta)$ with expected maximum coinciding with $\max(\mathbf{x})$.
If we frame this as a Bayesian argument we could say that we are tying to find a prior that aligns well with our observed data and our model for the data. To find a prior we consider a very simple family of discrete distributions based on the empirical distribution of our data. In the usual Bayesian arguments the data model is expressed as a likelihood, $f(\mathbf{x}\mid\theta)$. We rely instead on simulations for the more complex models above.
We could instead consider a Bayesian analysis where we specify a prior for $\theta$ and try to find a posterior distribution for $\max(\theta)$. Approximate Bayesian Computation (Rubin, 1984) is quite close to what we do above in that instead of a likelihood the analysis uses a simulation of the data generating process.
Specifically we could specify some weakly informative prior, $P(\theta)$, simulate data, $\mathbf{x}^{\prime}$, based on a draw, $\bar{\theta}\sim P(\theta)$, and check if $|\max(\mathbf{x}^{\prime})-\max(\mathbf{x})|<\epsilon$ for some acceptable $\epsilon>0$. If yes, we consider $\max(\bar{\theta})$ a draw from the posterior distribution, $P(\max(\theta)\mid\mathbf{x})$, otherwise we discard it. Repeating this would give us a posterior distribution for $\max(\theta)$ without us having to specify some ad-hoc family of priors, requiring instead a single prior distribution and the specification of $\epsilon$.
6 Discussion
Some of the insights presented here are obvious to anyone with a basic understanding of probability, like the introductory coin-flip example in Section 3.2. The mechanisms we have described are similar to those we find in regression to the mean, a well-known phenomenon in various fields such as livestock breeding and sport results. We believe that this has been neglected mainly because of the lack of formalisation: What is meant by SOTA performance, and what question does it answer? The main contribution has been to phrase the question, both the one that is commonly answered by reporting the maximum performance, and the question that can give an unbiased estimate of the SOTA performance.
Although our models span from the special case of iid classifiers to heterogeneous and correlated ones, the analysis is limited by the assumptions one has to make regarding distributions and correlation coefficients. An important lesson for practical applications is that simpler models tend to overestimate the bias, and hence underestimate the SOTA.
Looking at the application to real data in Section 5, we see that the implications of shifting the focus from best sample performance to best classifier performance has huge implications. Not necessarily so much in the new SOTA performance estimate, like the $0.005$ shift in Section 5.1.1, but definitely in which classifiers we consider performing at SOTA level - in this example, $25\%$ of the participating teams. In the Melanoma example in Section 5.2.3, the decrease in estimated SOTA performance is more striking, and we must keep in mind the wisdom from before, that simpler models tend to overestimate the bias. However, any model will uncover bias, and any bias will have quite dramatic results in terms of number of teams performing at SOTA level.
What does this mean for how we view the results from public challenges, such as Kaggle? First, we would like to emphasize that we appreciate the positive contributions that public challenges make, and that this is meant as a modification of how their results are viewed in the aftermath.
The beauty of the approach presented in this work - that we do not even have to know which classifier is the best to estimate the best performance - becomes the curse for the public challenge: We do not know which team is the best, only which was good combined with luck. The approach described here offers a method to quantify that luck, or uncertainty as we would name it in statistics. Like anyone with a basic understanding of probability can draw the conclusions in the coin-flip example, anyone with a basic understanding of statistics will look at the distribution of observed performances, as in Figure 8(a), and suspect that the right tail is a product of sample variance. This work offers a way of connecting the dots from coin-flip to public challenge, and reveals some of the plentiful hurdles along that route.
In some fields of science, where hundreds of thousands of tests are performed simultaneously, the effects of multiplicity simply cannot be ignored. But as we see from Figures 4(a) and 4(b), this is not a thousand-classifier problem, it is a two-classifier problem. As soon as you have two classifiers, the observed maximum is appreciably biased.
It is obvious from Figure 8(d) that the chosen approach to find a good empirical distribution is not optimal. A sample from the empirical distribution in Figure 8(c) should ideally result in a sample similar to the observed sample estimates, that is, an overlap between the black and the gray histograms in Figure 8(d). Whereas this is not that difficult to achieve, it requires more parameters and more tweaking, and does not result in any generalizable model. When deep-diving into a specific problem, it can be worthwhile spending some extra time refining those models, and presenting different results depending on model choice.
There are several approaches that can contribute to the problem of multiplicity, as presented in Section 2. Common for the solutions presented there are that they require additional data or more elaborate set-ups for public challenges. The approach presented here offers a tool that can be applied to already published results, and does not require convincing anyone else to change their preferred approach to finding the best solution.
Appendix A Numerical results
Expected values and standard deviations for $\operatorname{\hat{\theta}_{\max}(\mathbf{X})}$
[…]